OpenStudy (anonymous):

Is there a shortcut to quickly find a square of a number? let's say 65 squared..

OpenStudy (jtvatsim):

There are a lot of methods out there for "fast math tricks". If you Google "how to square numbers in your head" you will find a lot of info.

OpenStudy (anonymous):

Yes! (a + b)^2 = (a + b) x (a + b) = a^2 + 2ab + b^2 , where 60 is a ,and 5 is b. You can do this with any two digit number easily.

OpenStudy (jtvatsim):

Stole my idea @shawmoes ! :) (65)^2 = (60 + 5)^2 = 60^2 + 2*60*5 + 5^2 = 3600 + 2*300 + 25 = 4225.

hartnn (hartnn):

the 2 digit numbers ending in 5 are special cases :) 15^2 = 225 25^2 = 625 35^2 = 1225 45^2 = 2025 .... notice any pattern?? the last 2 digits are ALWAYS 25. the first part of the answer is alsways \(\large n \times (n+1)\) 85^2 \(\Large 8\times 9 =72\ \\\Large 85^2 = 7225\))

OpenStudy (mathmath333):

there is a good book for quick mental math with vast ideas for multiplication, addition , substraction, division. , i have read that and i think it very useful http://en.wikipedia.org/wiki/Trachtenberg_system \(\text{Trachtenberg system of speed mathmatics - Jakow Trachtenberg}\)